### v1

@@ Abstract
Conclusions: We provide insights into developing a framework for portable BCIs, bridging the gap between the laboratory and daily life. Specifically, our framework integrates modular design, real-time data processing, user-centred feedback, and low-cost hardware to deliver an accessible and adaptable BCI solution, addressing critical gaps in current BCI applications.
@@ 1 Introduction
Motivated by the Cybathlon 2024, we developed a modular, mobile EEG-based BCI system tailored to our pilot, but which can be easily adapted to other users. Our system was designed to address key challenges of portability, low latency, and user-centered adaptability while maintaining competitive decoding performance. The architecture integrates distinct modules for data acquisition, preprocessing, classification, and feedback, operating in both offline and online modes to optimize real-time performance. Despite rapid advances in BCI technology, achieving robust and reliable control in real-world environments, such as the dynamic and high-pressure setting of the Cybathlon, remains a major challenge. Leveraging recent advances in deep learning, we implement a classifier based on the linear recurrent neural network diagonalized Structured State-Space Sequence (S4D) layer to enhance classification accuracy and signal stability.
@@
    Robust online processing pipeline: Implementation of a low-latency, real-time architecture for EEG decoding, featuring artifact correction, adaptive re-referencing, and stable classification performance.
@@ 2.2 Task design
In addition to mental imagery task data, we also collect calibration data for artefact removal, which entails the participant sitting as relaxed and calm as possible, avoiding movement and eye blinks for one minute [Guarnieri2018].
Using this paradigm, we tested various mental tasks, including motor imagery, mental arithmetic, mental rotation, singing, and imagining the taste of a lemon. For mental arithmetic, the pilot was asked to multiply four-digit numbers shown on a sheet of paper, in order to avoid the task relying heavily on memory. Several combinations of tasks were also evaluated, since different combinations can influence performance, both for the user and the BCI classification [Kim2013].
@@ 2.4 Preprocessing
The pipeline begins by reading the raw EEG data files and corresponding event markers. It also loads one minute of calibration data and a mapping file that specifies which on-screen icons or marker labels correspond to the pilot’s specific actions or motor imagery tasks. The raw EEG data are then band-pass filtered between 1 and 40 Hz using a infinite impulse response (IIR) Chebyshev type 2 filter [Gramfort2021SelfSupervisionEEG], [Ouyang2025ProtocolEEG], [Delorme2004EEGLAB]. The filter has an almost constant group delay of 500ms. We forgo a notch filter since the bandpass filter sufficiently filters out the 50 Hz power line disturbance for our needs.
@@ 2.5 Feature Extraction
After the feature extraction, we perform a stratified 80/20 train-test split, maintaining prior data distributions. Finally, we apply windowing to extract more data points for training the classifier on.
@@ 2.6 Classifiers
Our BCI system supports the training and use of different machine learning and deep learning classifiers. We started with classical machine learning models. In particular, we utilized Support Vector Machines (SVM) extensively during the development process [Pedregosa2011ScikitLearn]. SVM operates by finding a hyperplane that best separates different classes in a high-dimensional space, optimizing the margin between classes for robust classification results [CortesVapnik1995SupportVectorNetworks]. This makes it suitable for generalizing to small datasets, such as ours. In addition to traditional approaches, we implemented a classifier consisting of stacks of bidirectionally running S4D layers [gu2022b] and a linear projection onto our set of classes. A state-space model, such as S4D, maps the history of an input sequence onto a set of basis polynomials [gu2020], [gu2022trainhippostatespace], which can be Legendre polynomials or sines and cosines [gu2022b], [gu2022efficientlymodelinglongsequences]. The history map corresponds to a linear recurrence, allowing for online updates of the hidden states [gu2020]. This approach is similar to classical signal processing techniques (e.g., the Fourier transform).
@@ 2.7.1 Games
The game features a QuickTime mode, in which the character pauses or slows down before an obstacle to create a temporally controlled interaction window. During this three-second phase, a cue marker appears, prompting the pilot to perform a specific motor imagery task. The classifier output is continuously monitored, and the character jumps only when the intended command is correctly detected above a predefined threshold. When misclassifications occur, a visual control bar decays in height, providing immediate and intuitive feedback on classifier performance.
@@ 3.3 Latency Analysis
Additional methodological latency stems from the group delay of the filter, which is $500$ ms, and from the windowing procedure, which introduces a further $1000$ ms. This results in a total delay of $1617$ ms, or a theoretical information transfer rate (ITR) [Billinger2013, SADEGHI2019101607] of $17.57$ bits per minute. We call this the theoretical ITR, since for online gameplay more than one classification is needed. A more realistic estimate should take into account the multiple classifications needed. From experimental observations, this can range from 1.51.533 seconds, resulting in an experimental ITR of 9.119.116.156.15 bits per minute.
@@ 4 Discussion
Our two main test pilots attempted various mental tasks: MI for the left and right hands, MI for the left and right halves of the body, MI for the arms and legs, mental singing, and mental multiplication. While one pilot was able to produce consistent MI for the legs, the second pilot was not able to do so. Hence, one chose the left-vs-right body half MI, including the legs, and the second only performed the hand MI. During the human-centered task design approach, mental multiplication was quickly abandoned because multiplying large numbers required significantly more time (up to ten seconds) than the three seconds chosen for the online EEG BCI paradigm, making it uncomfortable for the pilots. On the other hand, smaller numbers would have transformed the task into a purely memory-based retrieval. This demonstrates the importance of validating the pilot’s ability to perform specific mental tasks to produce activation, as well as the importance of tailoring BCIs to user capabilities.
@@
One main bottleneck we face with our system is delays. This reduces the usability of our system for daily use. Although the computational delay is comparatively low, the methodological delay is a significant issue. To have a BCI that can be used in closed-loop systems, delays below $250$ ms are recommended [sahar2016, xu2014, xu2016]. To reduce the delay, one could consider using deep learning frameworks that implicitly learn the filtering as one of their layers. This would remove the $500$ ms group delay from the Chebyshev filter. In principle, our S4D model could already perform this, since it can be interpreted as a convolution [gu2022efficientlymodelinglongsequences], [STUCHI2024111443]. This could help in cutting overall delay time. This would also remove any nonlinearities in the group delay, that might negatively affect the timeseries aspect of the S4D-layer based model through frequencies being assigned to a wrong point in time. However, since the used filter exhibited a relatively constant group delay, this was not an issue in our case.
@@
The S4D-layer-based model consistently achieves high accuracies and has demonstrated success in real-time use. This is accompanied by short training times of around 15 minutes on readily available high performance laptops. This is important, since training large models over multiple hours is not easily feasible, due to experimental time constraints and possible distribution shifts [RAZA2019154]. While the EEGEncoder achieved higher accuracies in offline testing, the longer training times could be a mitigating factor to make it a usable model for time-constrained experiments. Given the strong offline performance of the EEGEncoder architecture, future work should focus on assessing its real-time performance and integration into our BCI system. This could also include trading a few percentage points in accuracy for shorter training times. This highlights the importance of going beyond offline accuracies to real-world, real-time validation. Our results fall in line or outperform other systems in both offline [Lawhern2018EEGNet, Tabar2016DeepLearningEEG, 10.1371/journal.pone.0268880, 10.3389/fninf.2022.961089] and online [10.3389/fninf.2022.961089, signals4010004, JAFARIFARMAND2020101749] classification accuracies.
@@ 5 Conclusion
Our system exhibits low computational latency, with further improvements possible through adapted filtering and windowing procedures. We implement a deep learning classifier, the structured state-space sequence, which has not been previously used for EEG motor imagery classification. We achieve state-of-the-art classification performance while maintaining short training times. We also test an EEGEncoder architecture for EEG data published after the Cybathlon. While it outperformed the S4D-layer-based model, it also faced higher training times, paving the way for future optimization.
@@
Our findings underscore the importance of modularity and adaptability in BCI systems. By designing each module to function independently and in parallel, we make it easier and more feasible to substitute or upgrade components, such as preprocessing algorithms, classifiers, or feedback modalities, without compromising the overall system integrity. Furthermore, the system’s modularity aims to provide a promising foundation for scalable BCI solutions that can be personalized for different users, signal modalities, or assistive technologies.
@@
The datasets created for and subsequently used and analyzed during the current study are available from the corresponding author on reasonable request. The code used for this study will be made available upon reasonable request.

### v3

@@ Abstract
Conclusions: We provide insights into developing a framework for portable BCIs, taking a step towards bridging the gap between the laboratory and daily life. Specifically, our framework integrates modular design, real-time data processing, user-centred feedback, and low-cost hardware to deliver an accessible and adaptable BCI solution, addressing critical gaps in current BCI applications.
@@ 1 Introduction
Motivated by the Cybathlon 2024, we developed a modular, mobile EEG-based BCI system tailored to our pilot, but which can be easily adapted to other users. Our system was designed to address key challenges of portability, low latency, and user-centered adaptability while maintaining competitive decoding performance. The architecture integrates distinct modules for data acquisition, preprocessing, classification, and feedback, operating in both offline and online modes to optimize real-time performance. Despite rapid advances in BCI technology, achieving robust and reliable control in real-world environments, such as the dynamic and high-pressure setting of the Cybathlon, remains a major challenge. Leveraging recent advances in deep learning, we implement a classifier based on the linear recurrent neural network diagonalized Structured State-Space Sequence (S4D) layer to enhance classification accuracy and signal stability, which is novel to motor imagery classification.
@@
    Robust online processing pipeline: Implementation of a real-time architecture for EEG decoding, featuring artifact correction, adaptive re-referencing, and stable classification performance.
@@ 2.2 Task design
In addition to mental task data, we also collect calibration data for artefact removal, which entails the participant sitting as relaxed and calm as possible, avoiding movement and eye blinks for one minute [Guarnieri2018].
Using this paradigm, we also tested a mental arithmetic task. For mental arithmetic, the pilot was asked to multiply four-digit numbers shown on a sheet of paper, in order to avoid the task relying heavily on memory. Combinations of mental imagery and mental arithmetic tasks were also evaluated, since different combinations can influence performance, both for the user and the BCI classification [Kim2013].
@@ 2.4 Preprocessing
The pipeline begins by reading the raw EEG data files and corresponding event markers. It also loads one minute of calibration data and a mapping file that specifies which on-screen icons or marker labels correspond to the pilot’s specific actions or motor imagery tasks. The raw EEG data are then band-pass filtered between 4 and 36 Hz using an infinite impulse response (IIR) Chebyshev type 2 filter [Gramfort2021SelfSupervisionEEG], [Ouyang2025ProtocolEEG], [Delorme2004EEGLAB]. The filter has an almost constant group delay of 500ms. We forgo a notch filter since the bandpass filter sufficiently filters out the 50 Hz power line disturbance for our needs.
@@
The preprocessing is kept the same across models and only differs for the EEGEncoder architecture presented below (see section 2.6). Preprocessing parameters are presented in Table 1.
Table 1: Preprocessing parameters.
Method Parameter Value
Filter Frequency range 4-36 Hz
Filter Filter order 16
ASR Maximum dropout fraction 0.4
ASR Threshold 3 standard deviations
ASR Method Euclid
[…]
@@ 2.5 Feature Extraction
We perform a stratified 80/20 train-test split for all models, maintaining prior data distributions. Finally, we apply windowing to extract more data points for training the classifier on.
The feature extraction is performed the same way across all models, except the EEGEncoder, which is described below (see section 2.6). Feature extraction parameters are presented in Table 2.
Table 2: Feature extraction parameters.
Method Parameter Value
Morlet Number of cycles 2
Morlet Decimation factor 3
Morlet Feature frequencies 8-30 Hz
CSP Number of CSP components 6
[…]
@@ 2.6 Classifiers
Our BCI system supports the training and use of different machine learning and deep learning classifiers. We started with classical machine learning models. In particular, we utilized Support Vector Machines (SVM) extensively during the development process [Pedregosa2011ScikitLearn]. SVM operates by finding a hyperplane that best separates different classes in a high-dimensional space, optimizing the margin between classes for robust classification results [CortesVapnik1995SupportVectorNetworks]. This makes it suitable for generalizing to small datasets, such as ours. In addition to traditional approaches, we implemented a classifier consisting of three stacks of bidirectionally running S4D layers [gu2022b] and a linear projection onto our set of classes. A state-space model, such as S4D, maps the history of an input sequence onto a set of basis polynomials [gu2020], [gu2022trainhippostatespace], which can be Legendre polynomials or sines and cosines [gu2022b], [gu2022efficientlymodelinglongsequences]. The history map corresponds to a linear recurrence, enabling the hidden states to be updated online [gu2020]. Updating the hidden state allows the model to retain information from previous time steps, which informs the classification of a new input. This approach is similar to classical signal processing techniques (e.g., the Fourier transform).
The exact architecture is the following: first, the input features are mapped through a linear projection to hidden_dim dimensions. This then passes through num_layers of S4D blocks with skip connections and normalization. Afterwards, we perform temporal average pooling and perform logit classification for num_classes to have a probability as an output. Each S4D block consists of an S4D kernel that implements a diagonal state space model with matrices $A$ (diagonal with complex eigenvalues), $B$, $C$ (complex), and $D$ (skip connections), as well as the learnable discretization step size per channel, $dt$. The kernel is followed by a Gaussian error linear unit (GELU) activation function and a dropout layer with a dropout_rate. The output projection of the S4D block is a Conv1D layer followed by a gated linear unit (GLU). We provide our parameters in Table 3. The implementation of the model is optimized through Vandermonde multiplication to compute the discrete convolution kernel and FFT-based convolution.
Table 3: S4 classifier parameters.
Parameter Value
num_classes 3
hidden_dim 256
num_layers 3
dropout_rate 0.4
@@ 2.7.1 Games
The game features a QuickTime mode, in which the character pauses or slows down before an obstacle to create a temporally controlled interaction window. During this three-second phase, a cue marker appears, prompting the pilot to perform a specific motor imagery task.
The classifier output leads to an increase in the height of a bar visible on screen if the classification matches the cue marker. When misclassifications occur, a visual control bar decays in height, providing immediate and intuitive feedback on classifier performance. If the bar reaches a certain threshold within a predefined time (e.g., 3 seconds), then the character jumps over the obstacle. If the bar is below the threshold, no jump occurs. A jump is recorded as a successful task completion in online testing of the classifier, and no jump is consequently a failed task in the online assessment of the classifier.
@@ 3.3 Latency Analysis
Additional methodological latency stems from the group delay of the filter, which is $500$ ms, and from the windowing procedure, which introduces a further $1000$ ms. This results in a total delay of $1617$ ms, or a theoretical information transfer rate (ITR) [Billinger2013, SADEGHI2019101607] of $17.57$ bits per minute. We call this the theoretical ITR, since for online gameplay, more than one classification is needed. A more realistic estimate should take into account the multiple classifications needed. From experimental observations, this can range from 1.51.533 seconds, resulting in an experimental ITR of 9.119.116.156.15 bits per minute.
@@ 4 Discussion
Our two main test pilots attempted various mental tasks: MI for the left and right hands, MI for the left and right halves of the body, MI for the arms and legs, and mental multiplication. While one pilot was able to produce consistent MI for the legs, the second pilot was not able to do so. Hence, one chose the left-vs-right body half MI, including the legs, and the second only performed the hand MI. During the human-centered task design approach, mental multiplication was quickly abandoned because multiplying large numbers required significantly more time (up to ten seconds) than the three seconds chosen for the online EEG BCI paradigm, making it uncomfortable for the pilots. On the other hand, smaller numbers would have transformed the task into a purely memory-based retrieval. This demonstrates the importance of validating the pilot’s ability to perform specific mental tasks to produce activation, as well as the importance of tailoring BCIs to user capabilities.
The flexible structure of our pipeline allows the pilots to give feedback, which can immediately be integrated into the system. Therefore, the experimental sessions also consisted of regular discussions about the personal preferences of the pilots and the technical possibilities we have. This is what we believe to be human-centered design: One cannot expect a device to be used if the users are not directly involved in the production. In our case, this manifested through pilots implementing features in the games or in the pipeline themselves and testing them in their next experiment. Through this iterative and collaborative approach, we successfully integrated the diverse perspectives of patients and healthy participants into our pipeline.
We acknowledge that the number of participants (N = 2) is relatively small. Nevertheless, we believe that our system can be expanded to a wider user base, and we will strive to do so for future work. First, since our pipeline is, in principle, task agnostic, a pilot does not necessarily need to be able to produce motor imagery, as other tasks could be devised. Second, the modularity as well as the easy and fast parametrization enable us to quickly react to different user performances. A practical example: if certain classes are overclassified during online gameplay, the thresholds for how many classifications are needed to make the dino jump can be updated during the gameplay, without the need to restart. A current limitation we face is the need to retrain the classifier for every gameplay session. This results in an overhead of approximately 30 minutes before the system can be used online. This could deter new users from using the system. Combined with the need for a few sessions to get accustomed to the system and learn how to perform motor imagery effectively, we believe this might limit the interested user base. To mitigate this factor and improve our system, we are considering incorporating transfer learning methods based on Riemannian geometry in the future. This would enable us to train one large classifier that can be retrained for specific users with a smaller sample of data. This would also help for repeated sessions for a single user, as the same principle could be applied [riemanntransferYing], reducing the impact of possible data shifts [RAZA2019154].
@@
One main bottleneck we face with our system is delays. This reduces the usability of our system for daily use. Although the computational delay is comparatively low, the methodological delay is a significant issue. To have a BCI that can be used in closed-loop systems, delays below $250$ ms are recommended [sahar2016, xu2014, xu2016]. To reduce the delay, one could consider using deep learning frameworks that implicitly learn the filtering as one of their layers. This would remove the $500$ ms group delay from the Chebyshev filter. In principle, our S4D model could already perform this, since it can be interpreted as a convolution [gu2022efficientlymodelinglongsequences], [STUCHI2024111443]. This could help in cutting overall delay time. This would also remove any nonlinearities in the group delay, which might negatively affect the timeseries aspect of the S4D-layer-based model through frequencies being assigned to the wrong point in time. However, since the used filter exhibited a relatively constant group delay, this was not an issue in our case.
The delays also contribute to the comparatively low Cybathlon performance: simpler tasks, such as the moving wheelchair task, could be solved, as there was no constant movement. However, with more complex tasks, which require more precision, performance rapidly dropped, prompting us to skip these in the interest of time. Therefore, our system can already complete time-insensitive tasks, but struggles with more time-sensitive ones. This also highlights a limitation of testing BCIs in software: while games provide much leeway to ignore smaller errors, hardware applications mostly do not, without severely impacting user safety. For future work, we will thus strive to reduce latency to a minimum. Aside from latency, other factors, such as accuracy or false classification of other classes, can also reduce the effective control over the system. Improving those while maintaining the same latency would thus also improve overall system control.
The necessity of feature extraction methods was tested in an ablation study. We found that using only Morlet wavelets or CSP achieved respective accuracies of 82.9% and 66.6% on the dataset of the patient. This compares to an accuracy of 84.1% for the combined features. This improvement is relatively small, suggesting that CSP features might not be necessary. Future work will explore feature selection methods, such as the Fisher score [fisherscore], to further improve the number of features used.
@@
The S4D-layer-based model, which has not been used for motor imagery classification before, consistently achieves high accuracies and has demonstrated success in real-time use. This is accompanied by short training times of around 5 to 15 minutes on readily available high-performance laptops, depending on whether CUDA support is available. This is important, since training large models over multiple hours is not easily feasible, due to experimental time constraints and possible distribution shifts [RAZA2019154]. While the EEGEncoder achieved higher accuracies in offline testing, the longer training times (up to two hours) could be a mitigating factor, rendering it unusable for time-constrained experiments. Given the strong offline performance of the EEGEncoder architecture, future work should focus on assessing its real-time performance and integration into our BCI system. This could also include trading a few percentage points in accuracy for shorter training times. The machine learning models we tested required relatively short training periods (2-5 minutes) and are simple to implement, without the need for performant hardware. The EEGNet model falls somewhere in between the machine learning models and the S4D-layer-based model in terms of training times.
Recent MI-BCI studies employing transformer-based architectures have reported strong performance on standard benchmark datasets such as BCI Competition IV-2a and IV-2b. For example, MSCFormer achieved average accuracies of 82.60% on the four-class IV-2a dataset and 88.00% on the two-class IV-2b dataset [msconv], while CTNet reported subject-specific accuracies of 82.52% and 88.49% on IV-2a and IV-2b, respectively. In comparison, our approach achieves an average accuracy of 84% on a three-class motor imagery dataset [zhao2024ctnet]. Although these results are not directly comparable due to differences in class cardinality, dataset characteristics, and evaluation protocols, the obtained performance places our method within the range of recently reported state-of-the-art MI-BCI decoding approaches.
Based on this discussion, choosing a model becomes a trade-off between accuracy, training time, and feasibility to run the model on available hardware in online settings. This highlights the importance of going beyond offline accuracies to real-world, real-time validation. For our system, the S4D-layer-based model performed well, as its accuracy fell in line or outperformed other systems in both offline [Lawhern2018EEGNet, Tabar2016DeepLearningEEG, 10.1371/journal.pone.0268880, 10.3389/fninf.2022.961089] and online [10.3389/fninf.2022.961089, signals4010004, JAFARIFARMAND2020101749] classification settings. The training times did not negatively affect experiment execution, and we deem our hardware sufficient and accessible, making it a good option for our use case.
@@ 5 Conclusion
Our system exhibits low computational latency, with further improvements possible through adapted filtering and windowing procedures. We implement a deep learning classifier, the structured state-space sequence, which has not been previously used for EEG motor imagery classification. We achieve sufficient classification performance while maintaining short training times. We also test an EEGEncoder architecture for EEG data published after the Cybathlon. While it outperformed the S4D-layer-based model, it also faced higher training times, paving the way for future optimization.
@@
Our findings underscore the importance of modularity and adaptability in BCI systems. By designing each module to function independently and in parallel, we make it easier and more feasible to substitute or upgrade components, such as preprocessing algorithms, classifiers, or feedback modalities, without compromising the overall system integrity. Furthermore, the system’s modularity could provide a foundation for more scalable BCI solutions that can be personalized for different users, signal modalities, or assistive technologies.
@@
  • GLU: gated linear unit
  • GELU: Gaussian error linear unit
@@
The datasets created for and subsequently used and analyzed during the current study are available under [neurotum_e_v_2025_18087806]. The pilot IDs are P1 for the patient and P2 for the healthy participant. The code used for this study will be made available upon reasonable request.